OpenStudy (anonymous):

Solve the given equation for x. 8x-7-4x=x+4+x Select the correct choice below and, if necessary, fill in the answer box to complete your choice. A. x=? B. The solution is all real numbers. C. There is no solution

Nnesha (nnesha):

first combine like terms on left and right side

Nnesha (nnesha):

8x -4x= ??? x + x =??? there is one at front of x

OpenStudy (anonymous):

4x

Nnesha (nnesha):

yes so 4x on left side what about x+x

OpenStudy (anonymous):

2x

Nnesha (nnesha):

yes right so now you have 4x - 7 = 2x + 4 first add 7 both side

Nnesha (nnesha):

 ??